OpenStudy (anonymous):

Kevin Sold A house for $57000. His fee or sales commission for selling the house was $2679. What percent of the price of the house was kevins commission?

hero (hero):

is/of = %/100 2679/57000 = x/100 267900 = 57000x 267900/57000 = x 2679/570 =x 4.7 = x Therefore, 2679 is 4.7% of 57000